OpenStudy (anonymous):

what is the sum of this infinite geometric series? 2+ 2/5+2/25+2/125+.....

OpenStudy (salina):

ratio = (2/5)/2=1/5

OpenStudy (anonymous):

\[\sum \frac{1}{5^k}=\frac{1}{1-\frac{1}{5}}\]

OpenStudy (salina):

you have 4 term

OpenStudy (anonymous):

get \[2\times \frac{1}{1-\frac{1}{5}}=2\times \frac{1}{\frac{4}{5}}=2\times \frac{5}{4}=\frac{5}{2}\]

OpenStudy (anonymous):

a. sideway 8 b.2.5 c.1.25 d.10 e.5

OpenStudy (anonymous):

5/2 = 2.5 last time I checked.